# Contestable AI needs Computational Argumentation

### v1

@@ 1 Introduction
AI has become pervasive in recent years, with applications ranging from autonomous driving (?) to finance (?) and healthcare (?). Current state-of-the-art AI systems focus on algorithmic solutions, often built from data, generating outputs in the forms of predictions, recommendations and/or decisions and, in some cases, explanations thereof. These existing solutions, though, mostly neglect the need for these AI systems to be contestable. Instead, the need to accommodate contestability is a crucially important problem if AI systems are to be deployed in and benefit society. Indeed, contestability is prominently advocated in some frameworks for AI ethics and in regulations. For example, the Organisation for Economic Co-operation and Development (OECD) states that information should be provided “to enable those adversely affected by an AI system to challenge its outcome based on plain and easy-to-understand information on the factors, and the logic that served as the basis for the prediction, recommendation or decision” (Principle 1.3111https://oecd.ai/en/dashboards/ai-principles/P7.). Moreover, the ACM Global Technology Policy Council for Responsible Algorithmic Systems advocates ‘Contestability and Auditability’ as one of nine principles.222https://www.acm.org/media-center/2022/october/tpc-statement-responsible-algorithmic-systems. Contestability is required by law in some jurisdictions, e.g. GDPR, article 22(3),333https://gdpr-text.com/read/article-22/. states that the data subject’s rights to be safeguarded shall include “at least the right […] to contest the decision”. Finally, contestability and redress are also identified as key principles underpinning the UK AI regulation framework 444https://www.gov.uk/government/publications/ai-regulation-a-pro-innovation-approach/white-paper.
@@
Contestability is seen by some (?) as a means to facilitate accountability, e.g. to prevent the use of Uber’s algorithm for banning drivers because it is not taking into account bias of customers when giving bad reviews. A handful of recent works (overviewed in Section 2) bring contestability to the attention of AI researchers, practitioners and users, while providing insights into its possible interpretations in practice. Moreover, there is evidence, in the form of user studies, that contestability can affect users’ perception of AI fairness (?). However, formal/computational forms of contestability are mostly lacking in the literature, with some exceptions, in the form of 1. indications that, in the case of machine learning, methods generating counterfactual explanations in terms of “actionable changes that an individual can make to flip the prediction of the classifier” may be interpreted as offering (limited) contestability (?) and 2. (?), offering contestability in the very specific setting of causal discovery. Finally, existing works mostly see contestability as a post-hoc process, after predictions/recommendations/decisions have been computed.
@@
In this position paper (specifically in Section 3) we consider what it means for (any) algorithmic decision systems (ADS) (?) to be contestable and contested (by humans or other ADSs). Our take-away message is that formalising/realising computationally contestable AI requires: explanations (by the ADSs which humans/other ADSs may want to contest); grounds for contestation (by the contesters); ability (by the contested) to redress any issues successfully raised during contestation; and ability to interact (by both contested and contester). We then advocate computational argumentation (CA) as being ideally suited to support contestability, providing evidence from the literature on CA (Section 4). We finally conclude (Section 5) with some pointers for future directions in what we believe is a very promising direction of research for the KR community.
@@ 2 AI Contestability in the literature
Contested entities. Prior work has focused on contesting various aspects of an ADS, including its general design (?; ?; ?), training data (?), training procedure (?), its inputs (?) and outputs (?; ?; ?; ?; ?; ?) as well as the ADS as a whole (?; ?; ?; ?; ?; ?). Note that the possibility of humans being contested by an ADS (e.g., if their assumptions about the subject domain are flawed) is not typically considered.
@@
Contesting entities. All approaches we identified focus on humans as the contesting entities. However, there are considerable differences between the categories and assumed skill sets of people engaged in contestation. Several argue that decision subjects should be empowered to directly contest ADS decisions affecting them, potentially without possessing detailed knowledge of the subject domain (?; ?; ?; ?; ?; ?). Some works suggest that individual decision subjects may not always be able to effectively contest on their own, and propose contestation via third-party representatives (?) or as part of a group in a “class action” (?). Others also consider contestation by professionals, subject matter experts and regulators (?; ?; ?; ?; ?). Note that the possibility of ADSs being the contesting entities is not typically considered.
@@
Contestation methods. Several methods have been proposed for facilitating contestation. In the realm of non-technical solutions, participatory design has been advocated as a way to involve various stakeholders in ADS development, enabling advance mitigation of possible issues and risks associated with the use of the ADS (?; ?; ?). The framework proposed in (?) suggests following a set of development practices over the ADS lifecycle, including incorporating ex-ante safeguards, gathering end-user feedback, implementing quality assurance, mitigating possible risks and allowing for third-party oversight. More technical approaches envision usage of operation logs (?) or model explanations (?; ?; ?; ?; ?). We discuss these and their use of explanations in greater detail next.
@@
The role of explanations. That explanations are needed to support contestability is acknowledged by some (e.g. by (?; ?; ?; ?)), but the problem of using explanations to support contestability in practice has received little attention in explainable AI (XAI) (or AI for that matter) to date. A notable exception is (?), where causal discovery with neural networks guides human feedback for contesting the discovered causal relations, which can be seen as a form of global explanation. Also, some current XAI methods can be seen as offering ground for some limited contestability when they suggest actionable recourse, as is the case for counterfactual explanations in terms of “actionable changes that an individual can make to flip the prediction of the classifier” (?). However, these works disregard that algorithmic decisions may actually be incorrect, e.g. because they are based on incorrect or incomplete data. Moreover, these methods are one-shot, providing no opportunities for follow-up inquiry, and shallow, revealing no information about the steps or logic that led to the explained output, and thus offer little ground for contestability. Some other XAI methods provide information on the AIs’ deliberation and insecurities (?); however, they do not support any form of contestability. Overall, explanations, as understood in state-of-the-art XAI, are seen as inadequate to support contestability (?). Specifically, (?) advocate a generic notion of explanations that capture the rationales behind the development and deployment of the ADS (referred to therein as “process-centric explanations”).
@@
Post-hoc and ex-ante contestability. Some take the view that contestability needs to be supported by regulatory frameworks, involving in particular policy modelling and normative reasoning (?). These approaches see contestability as post-hoc processes, detached from explainability, “to review algorithmic decisions” (?). Ex-ante contestability is envisaged by some, e.g. (?), as a design principle (?), whereby systems are designed to enable users to interactively contest these systems and interaction between humans and systems is needed for “critique and correction” (?), but no technical solutions exist on how to support it.
@@
Contestability as an interaction process. (?) see contestability as an interaction process with humans, at development and deployment times, to allow humans, rather than just data, to train systems. Several other works (e.g. (?; ?)) envisage post-hoc contestability as an interaction process, again between humans and ADSs alone.
@@
Other related work. Works on corrigibility, such as by (?; ?; ?) focus on human overseers providing feedback/instructions to ADSs so that they align with the intentions/values of their user(s) (?; ?). While corrigibility can be viewed as enabling human users to contest the decisions of ADSs, it is narrowly focused on ADSs which optimise for reward when they have incentives to manipulate users (?), ignore instructions (?; ?), and disempower humans/other ADSs (?).
@@ 3 An abstract view of AI contestability
Explanations for Contestability. In line with some of the literature, we envisage that contestability needs to be supported by explanations for and/or its outputs for specific inputs. We thus assume that the ADS, in addition to , is equipped with an explanation method . Differently from (?), we believe that technical notions of explanations from the literature in XAI may already prove useful. These explanations may be local (i.e. relating to a specific input, output pair) or global (i.e. relating to all input, output pairs). Different forms of contestability may require and/or benefit from different forms of explanations and thus explanation methods, including as follows: 1. a feature-attribution explanation method, e.g. as provided by model-agnostic LIME (?) or SHAP (?) or by model-specific (tailored to neural classifiers) LRP (?) or DeepLIFT (?); 2. an abductive explanation method, e.g. as in (?); 3. a counterfactual explanation method e.g. as first envisaged in (?; ?); 4. a rule-based explanation method e.g as in (?; ?; ?) 5. a mechanistic (local) explanation method, e.g. as provided in DAX (?), SpArX (?), and (?). 6. a mechanistic (global) explanation e.g as in (?); 7. a surrogate model for interpreting , e.g. (?; ?; ?; ?); 8. itself, if already an interpretable model (e.g. a decision tree, or a set of rules).
We stress that satisfying some desirable properties will make the explanation methods more useful to the contestation. Notably, faithfulness requires that the explanations reflect the true reasoning of the explained models (?). This will guide the interactions in contestations towards the correct direction, and allow for more effective redress. Another important property is robustness (?). Non-robust explanation methods could result in drastically different explanations for two users with similar inputs (?), which may jeopardise their explanatory function. In contrast, robust methods (?; ?; ?; ?) improve the consistency and trustworthiness of explanations and may be better suited to support contestability.
@@
Interaction for contestability. The process of contestation needs to be supported by suitable forms of interaction between the contested ADS and the (human or ADS) contester. Interaction may be in the form of conversations in natural language, e.g. using explanations generated by Large Language Models (?), especially if the contester is a human. Alternatively, they may be in structured format, using a formal agent communication language, e.g. in the spirit of FIPA (?).999http://www.fipa.org/. In Figure 1, we indicate with the method used by contested ADS and contester to engage in the interactions necessary for contestability.
@@
Redress. While explanations empower contestation, its complete realisation needs the ADS to have the ability to redress any issues (successfully) brought about in the contestation. Thus, we assume that the ADS is also equipped with a redress method . For illustration, in the loan application Example 1, may amount to revising the first rule by allowing for exceptions to be made when the breaks are due to training or parental leaves. As an additional illustration for text classifiers, redress may result from the a post-hoc reasoning process with the classifier’s outputs and external knowledge encoded in argument schemes, as in (?), or with the classifier’s outputs and explanations therefor, e.g. as in (?).
Note that we see contestation, and thus redress, as a post-training process, rather than model debugging during training, such as via data augmentation (e.g. as in (?)) or regularisation (e.g. as in (?; ?; ?; ?)). Thus, our view is that contestation is, in general, different from explanation-based model debugging (e.g. as in (?) for tabular data, (?) for text classification, or (?) for image classification). We, however, envisage that, in the case of models trained from data, redress may at times involve fine-tuning or retraining steps in which case the above methods, as well as methods for repairing AI models, e.g. (?; ?; ?), could be drawn upon. In the illustrative case of a black-box text classifier, an example of contestation of the full model (the third setting) is offered by FIND (?): here, LRP (?) is used to associate output neurons in the feature extractor of a text classifier with word clouds. Users can then contest the use of individual neurons by disabling them and fine-tuning the model to no longer rely on them. This allows, specifically, to decrease model bias and reliance on artifacts (?). (?), similarly, aim to expose visual concepts learned by an image classifier using concept activation vectors (?) and fine-tune the model to mitigate the use of specific concepts using gradient regularisation.

### v2

@@ 1 Introduction
AI has become pervasive in recent years, with applications ranging from autonomous driving (Muhammad et al., 2020) to finance (Cao, 2022) and healthcare (Shaheen, 2021). Current state-of-the-art AI systems focus on algorithmic solutions, often built from data, generating outputs in the forms of predictions, recommendations and/or decisions and, in some cases, explanations thereof. These existing solutions, though, mostly neglect the need for these AI systems to be contestable. Instead, the need to accommodate contestability is a crucially important problem if AI systems are to be deployed in and benefit society. Indeed, contestability is prominently advocated in some frameworks for AI ethics and in regulations. For example, the Organisation for Economic Co-operation and Development (OECD) states that information should be provided “to enable those adversely affected by an AI system to challenge its outcome based on plain and easy-to-understand information on the factors, and the logic that served as the basis for the prediction, recommendation or decision” (Principle 1.3111https://oecd.ai/en/dashboards/ai-principles/P7). Moreover, the ACM Global Technology Policy Council for Responsible Algorithmic Systems advocates ‘Contestability and Auditability’ as one of nine principles.222https://www.acm.org/media-center/2022/october/tpc-statement-responsible-algorithmic-systems Contestability is required by law in some jurisdictions, e.g. GDPR, article 22(3),333https://gdpr-text.com/read/article-22/ states that the data subject’s rights to be safeguarded shall include “at least the right […] to contest the decision”. Finally, contestability and redress are also identified as key principles underpinning the UK AI regulation framework444https://www.gov.uk/government/publications/ai-regulation-a-pro-innovation-approach/white-paper, whereby “Guidance should clarify existing ‘formal’ routes of redress offered by regulators in certain scenarios”. This is a major departure point from “classical” requirements for explainable systems, which typically do not require correcting the AI when faulty behaviours are exposed.
@@
Contestability is seen by some (Hicks, 2022) as a means to facilitate accountability, e.g. to prevent the use of Uber’s algorithm for banning drivers because it is not taking into account bias of customers when giving bad reviews. A handful of recent works (overviewed in Section 2) bring contestability to the attention of AI researchers, practitioners and users, while providing insights into its possible interpretations in practice. Moreover, there is evidence, in the form of user studies, that contestability can affect users’ perception of AI fairness (Yurrita et al., 2023). However, formal/computational forms of contestability are mostly lacking in the literature, with some exceptions, in the form of 1. indications that, in the case of machine learning, methods generating counterfactual explanations in terms of “actionable changes that an individual can make to flip the prediction of the classifier” may be interpreted as offering (limited) contestability (Venkatasubramanian and Alfano, 2020) and 2. (Russo and Toni, 2023), offering contestability in the very specific setting of causal discovery. Finally, existing works mostly see contestability as a post-hoc process, after predictions/recommendations/decisions have been computed.
@@
In this position paper (specifically in Section 3) we consider what it means for (any) algorithmic decision systems (ADS) (Henin and Métayer, 2022) to be contestable and contested (by humans or other ADSs). Our take-away message is that formalising/realising computationally contestable AI requires: explanations (by the ADSs which humans/other ADSs may want to contest); grounds for contestation (by the contesters); ability (by the contested) to redress any issues successfully raised during contestation; and ability to interact (by both contested and contester). We then advocate computational argumentation (CA) as being ideally suited to support contestability, providing evidence from the literature on CA (Section 4). We finally conclude (Section 5) with some pointers for future directions in what we believe is a very promising direction of research for the KR community.
@@ 2 AI Contestability in the literature
Contested entities. Prior work has focused on contesting various aspects of an ADS, including its general design (Almada, 2019; Alfrink et al., 2022; Yurrita, Balayn, and Gadiraju, 2023), training data (Kluttz, Kohli, and Mulligan, 2022), training procedure (Almada, 2019), its inputs (Alfrink et al., 2022) and outputs (Hirsch et al., 2017; Ploug and Holm, 2020; Tubella et al., 2020; Lyons, Velloso, and Miller, 2021; Henin and Métayer, 2022; Hicks, 2022) as well as the ADS as a whole (Lyons, Velloso, and Miller, 2021; Vaccaro et al., 2021; Henin and Métayer, 2022; Hicks, 2022; Alfrink et al., 2023; Russo and Toni, 2023). Note that the possibility of humans being contested by an ADS (e.g., if their assumptions about the subject domain are flawed) is not typically considered.
@@
Contesting entities. All approaches we identified focus on humans as the contesting entities. However, there are considerable differences between the categories and assumed skill sets of people engaged in contestation. Several argue that decision subjects should be empowered to directly contest ADS decisions affecting them, potentially without possessing detailed knowledge of the subject domain (Almada, 2019; Ploug and Holm, 2020; Lyons, Velloso, and Miller, 2021; Vaccaro et al., 2021; Henin and Métayer, 2022; Hicks, 2022). Some works suggest that individual decision subjects may not always be able to effectively contest on their own, and propose contestation via third-party representatives (Alfrink et al., 2022) or as part of a group in a “class action” (Lyons, Velloso, and Miller, 2021). Others also consider contestation by professionals, subject matter experts and regulators (Tubella et al., 2020; Henin and Métayer, 2022; Kluttz, Kohli, and Mulligan, 2022; Alfrink et al., 2023; Russo and Toni, 2023). Note that the possibility of ADSs being the contesting entities is not typically considered.
@@
Contestation methods. Several methods have been proposed for facilitating contestation. In the realm of non-technical solutions, participatory design has been advocated as a way to involve various stakeholders in ADS development, enabling advance mitigation of possible issues and risks associated with the use of the ADS (Almada, 2019; Vaccaro et al., 2021; Alfrink et al., 2023). The framework proposed in (Alfrink et al., 2022) suggests following a set of development practices over the ADS lifecycle, including incorporating ex-ante safeguards, gathering end-user feedback, implementing quality assurance, mitigating possible risks and allowing for third-party oversight. More technical approaches envision usage of operation logs (Hicks, 2022) or model explanations (Almada, 2019; Ploug and Holm, 2020; Kluttz, Kohli, and Mulligan, 2022; Russo and Toni, 2023; Yurrita, Balayn, and Gadiraju, 2023). We discuss these and their use of explanations in greater detail next.
@@
The role of explanations. That explanations are needed to support contestability is acknowledged by some (e.g. by Almada (2019); Lyons, Velloso, and Miller (2021); Alfrink et al. (2022); Wachter, Mittelstadt, and Russell (2017)), but the problem of using explanations to support contestability in practice has received little attention in explainable AI (XAI) (or AI for that matter) to date. A notable exception is (Russo and Toni, 2023), where causal discovery with neural networks guides human feedback for contesting the discovered causal relations, which can be seen as a form of global explanation. Also, some current XAI methods can be seen as offering ground for some limited contestability when they suggest actionable recourse, as is the case for counterfactual explanations in terms of “actionable changes that an individual can make to flip the prediction of the classifier” (Venkatasubramanian and Alfano, 2020). However, these works disregard that algorithmic decisions may actually be incorrect, e.g. because they are based on incorrect or incomplete data. Moreover, these methods are one-shot, providing no opportunities for follow-up inquiry, and shallow, revealing no information about the steps or logic that led to the explained output, and thus offer little ground for contestability. Some other XAI methods provide information on the AIs’ deliberation and insecurities (Wang and Vasconcelos, 2019); however, they do not support any form of contestability. Overall, explanations, as understood in state-of-the-art XAI, are seen as inadequate to support contestability (Henin and Métayer, 2022). Specifically, Yurrita, Balayn, and Gadiraju (2023) advocate a generic notion of explanations that capture the rationales behind the development and deployment of the ADS (referred to therein as “process-centric explanations”).
@@
Post-hoc and ex-ante contestability. Some take the view that contestability needs to be supported by regulatory frameworks, involving in particular policy modelling and normative reasoning (Tubella et al., 2020). These approaches see contestability as post-hoc processes, detached from explainability, “to review algorithmic decisions” (Lyons, Velloso, and Miller, 2021). Ex-ante contestability is envisaged by some as a design principle (Hirsch et al., 2017; Alfrink et al., 2022), whereby systems are designed to enable users to interactively contest these systems and interaction between humans and systems is needed for “critique and correction” (Lyons, Velloso, and Miller, 2021), but no technical solutions exist on how to support it.
@@
Contestability as an interaction process. Kluttz, Kohli, and Mulligan (2022) see contestability as an interaction process with humans, at development and deployment times, to allow humans, rather than just data, to train systems. Several other works (e.g. (Hirsch et al., 2017; Lyons, Velloso, and Miller, 2021)) envisage post-hoc contestability as an interaction process, again between humans and ADSs alone.
@@
Other related work. Works on corrigibility, such as (Soares et al., 2015; Carey and Everitt, 2023; Russell, Dewey, and Tegmark, 2015) focus on human overseers providing feedback/instructions to ADSs so that they align with the intentions/values of their user(s) (Gabriel, 2020; Hadfield-Menell et al., 2016). While corrigibility can be viewed as enabling human users to contest the decisions of ADSs, it is narrowly focused on ADSs which optimise for reward when they have incentives to manipulate users (Ward, Toni, and Belardinelli, 2022), ignore instructions (Hadfield-Menell et al., 2016), and disempower humans/other ADSs (Turner et al., 2021).
@@ 3 An abstract view of AI contestability
Explanations for contestability. In line with some of the literature, we envisage that contestability needs to be supported by explanations for and/or its outputs for specific inputs. We thus assume that the ADS, in addition to , is equipped with an explanation method . Differently from Yurrita, Balayn, and Gadiraju (2023), we believe that technical notions of explanations from the literature in XAI may already prove useful. These explanations may be local (i.e. relating to a specific input, output pair) or global (i.e. relating to all input, output pairs). Different forms of contestability may require and/or benefit from different forms of explanations and thus explanation methods, including as follows: 1. a feature-attribution explanation method, e.g. as provided by model-agnostic LIME (Ribeiro, Singh, and Guestrin, 2016) or SHAP (Lundberg and Lee, 2017) or by model-specific (tailored to neural classifiers) LRP (Montavon et al., 2019) or DeepLIFT (Shrikumar, Greenside, and Kundaje, 2017); 2. an abductive explanation method, e.g. as in (Ignatiev, Narodytska, and Marques-Silva, 2019); 3. a counterfactual explanation method e.g. as first envisaged in (Wachter, Mittelstadt, and Russell, 2017; Tolomei et al., 2017); 4. a rule-based explanation method e.g as in (Ribeiro, Singh, and Guestrin, 2018; Guidotti et al., 2019; Ferreira et al., 2022) 5. a mechanistic (local) explanation method, e.g. as provided in DAX (Albini et al., 2020), SpArX (Ayoobi, Potyka, and Toni, 2023), and (Wang et al., 2023). 6. a mechanistic (global) explanation e.g as in (Ayoobi, Potyka, and Toni, 2023); 7. a surrogate model for interpreting , e.g. (Kenny and Keane, 2019; Tan and Kotthaus, 2022; Potyka, Yin, and Toni, 2023); 8. itself, if already an interpretable model (e.g. a decision tree, or a set of rules).
We stress that satisfying some desirable properties will make the explanation methods more useful to the contestation. Notably, faithfulness requires that the explanations reflect the true reasoning of the explained models (Lakkaraju et al., 2019). This will guide the interactions in contestations towards the correct direction, and allow for more effective redress. Another important property is robustness (Jiang et al., 2024b). Non-robust explanation methods could result in drastically different explanations for two users with similar inputs (Leofante and Potyka, 2024), which may jeopardise their explanatory function. In contrast, robust methods (Jiang et al., 2023; Leofante and Lomuscio, 2023; Leofante, Botoeva, and Rajani, 2023; Jiang et al., 2024a) improve the consistency and trustworthiness of explanations and may be better suited to support contestability.
@@
Interaction for contestability. The process of contestation needs to be supported by suitable forms of interaction between the contested ADS and the (human or ADS) contester. Interaction may be in the form of conversations in natural language, e.g. using explanations generated by Large Language Models (Bills et al., 2023), especially if the contester is a human. Alternatively, they may be in structured format, using a formal agent communication language, e.g. in the spirit of FIPA (Poslad, 2007).999http://www.fipa.org/ In Figure 1, we indicate with the method used by contested ADS and contester to engage in the interactions necessary for contestability.
@@
Redress. While explanations empower contestation, its complete realisation needs the ADS to have the ability to redress any issues (successfully) brought about in the contestation. Thus, we assume that the ADS is also equipped with a redress method . For illustration, in the loan application Example 1, may amount to revising the first rule by allowing for exceptions to be made when the breaks are due to training or parental leaves. As an additional illustration for text classifiers, redress may result from the a post-hoc reasoning process with the classifier’s outputs and external knowledge encoded in argument schemes, as in (Carstens and Toni, 2017), or with the classifier’s outputs and explanations therefor, e.g. as in (Freedman et al., 2024).
Note that we see contestation, and thus redress, as a post-training process, rather than model debugging during training, such as via data augmentation (e.g. as in (Teso and Kersting, 2019)) or regularisation (e.g. as in (Ross, Hughes, and Doshi-Velez, 2017; Rieger et al., 2020; Shao et al., 2021; Zhang, Williams, and Toni, 2024)). Thus, our view is that contestation is, in general, different from explanation-based model debugging (e.g. as in (Ghai et al., 2021) for tabular data, (Lertvittayakumjorn and Toni, 2021) for text classification, or (Popordanoska, Kumar, and Teso, 2020) for image classification). We, however, envisage that, in the case of models trained from data, redress may at times involve fine-tuning or retraining steps in which case the above methods, as well as methods for repairing AI models, e.g. (Henriksen, Leofante, and Lomuscio, 2022; Almog and Kalech, 2023), could be drawn upon. In the illustrative case of a black-box text classifier, an example of contestation of the full model (the third setting) is offered by FIND (Lertvittayakumjorn, Specia, and Toni, 2020): here, LRP (Montavon et al., 2019) is used to associate output neurons in the feature extractor of a text classifier with word clouds. Users can then contest the use of individual neurons by disabling them and fine-tuning the model to no longer rely on them. This allows, specifically, to decrease model bias and reliance on artifacts (Lertvittayakumjorn, Specia, and Toni, 2020). Dreyer et al. (2024), similarly, aim to expose visual concepts learned by an image classifier using concept activation vectors (Kim et al., 2018) and fine-tune the model to mitigate the use of specific concepts using gradient regularisation.